# BF-STVSR: B-Splines and Fourier—Best Friends for High Fidelity Spatial-Temporal Video Super-Resolution

### v1

@@ 5 Conclusions
In this paper, we proposed BF-STVSR, a novel framework for Continuous Spatial-Temporal Video Super Resolution (C-STVSR). We introduced two specialized position encoding modules: B-spline Mapper, which leverages B-spline basis functions for smooth and accurate temporal interpolation, and Fourier Mapper, which captures dominant spatial frequencies to effectively model fine-grained spatial details. Our analysis highlights that naive position encoding can degrade performance, particularly when paired with optical flow networks, emphasizing the importance of tailored encodings for spatial and temporal axes. Through extensive experiments, we show that BF-STVSR achieves state-of-the-art performance in PSNR and SSIM metrics across diverse datasets, demonstrating enhanced spatial detail, natural temporal consistency, and robustness under challenging conditions, including extreme out-of-distribution scales.

### v2

@@ 5 Conclusions
In this paper, we proposed BF-STVSR, a novel framework for Continuous Spatial-Temporal Video Super-Resolution (C-STVSR). Motivated by our observation that naïve position encoding can degrade performance---particularly when paired with optical flow networks---we introduced two axis-specific position encoding modules: B-spline Mapper, which leverages B-spline basis functions for smooth and accurate temporal interpolation, and Fourier Mapper, which captures dominant spatial frequencies to effectively model fine-grained spatial details. By estimating motion directly from encoded features, our design eliminates the need for external optical flow supervision, achieving high efficiency while maintaining strong performance. Extensive experiments confirm that BF-STVSR achieves state-of-the-art results in PSNR, SSIM and various video quality metrics, demonstrating superior spatial detail, natural temporal consistency, and robustness under challenging conditions, including extreme out-of-distribution scales.
Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.2022R1C1C100849612) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2020-II201336, Artificial Intelligence Graduate School Program (UNIST), No.2022-0-00959, No.RS-2022-II220959 (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making, RS-2022-II220264, Comprehensive Video Understanding and Generation with Knowledge-based Deep Logic Neural Network), and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City. This research used high performance computing resources of the UNIST Supercomputing Center.